# Calibrated Probabilistic Interpolation for GEDI Biomass

### v2

@@ Abstract
Reliable wall-to-wall biomass mapping from NASA’s GEDI mission requires interpolating sparse LIDAR observations across heterogeneous landscapes. While machine learning approaches like Random Forest and XGBoost are standard for this task, they treat spatial predictions of GEDI observations from multispectral or SAR remote sensing data as independent without adapting to the varying difficulty of heterogeneous landscapes. We demonstrate these approaches generally fail to produce calibrated prediction intervals. We identify that this stems from conflating ensemble variance with aleatoric uncertainty and ignoring local spatial context.
To resolve this, we introduce Attentive Neural Processes (ANPs), a probabilistic meta-learning architecture that explicitly conditions predictions on local observation sets and geospatial foundation model embeddings. Unlike static ensembles, ANPs learn a flexible spatial covariance function, allowing uncertainty estimates to expand in complex landscapes and contract in homogeneous areas. We validate this approach across five distinct biomes ranging from tropical Amazonian forests to boreal, temperate, and alpine ecosystems, demonstrating that ANPs achieve competitive accuracy while maintaining near-ideal uncertainty calibration. We demonstrate the operational utility of the method through few-shot adaptation, where the model recovers most of the performance gap in cross-region transfer using minimal local data. This work provides a scalable, theoretically rigorous alternative to ensemble variance for continental scale earth observation.
@@
{highlights}
Current GEDI interpolations derived from sparse GEDI LIDAR data rely on tree-based ensembles (Random Forest, XGBoost) that produce miscalibrated uncertainty estimates.
Attentive Neural Processes (ANPs) achieve competitive biomass estimation accuracy while maintaining near-ideal probabilistic calibration.
Cross-regional validation across biomes demonstrates that ANPs exhibit robust uncertainty expansion under spatial extrapolation.
The method enables scalable, context-aware uncertainty quantification for large-scale mapping.
@@ 1 Introduction
Forest aboveground biomass density (AGBD) is important for carbon accounting, climate change mitigation, and conservation planning (Dubayah et al., 2022; Duncanson et al., 2022). Accurate and high resolution biomass maps enable monitoring of deforestation, forest degradation, and carbon sequestration which are essential for implementing conservation programs and climate targets. NASA’s Global Ecosystem Dynamics Investigation (GEDI) mission provides high quality lidar measurements of forest structure, including calibrated estimates of AGBD at 25m footprint resolution (Dubayah et al., 2022). However, GEDI samples sparsely, with footprints separated by approximately 600m along-track in criss-crossing orbital patterns. To support operational forest monitoring, these sparse observations are interpolated to create wall-to-wall maps.
@@
Existing approaches to wall-to-wall biomass mapping from GEDI data typically aggregate to coarse resolution (1km) to increase sampling density (Duncanson et al., 2022; Dubayah et al., 2022). Recently, machine learning and deep learning methods have enabled finer-resolution mapping (10-30m) by combining GEDI footprints with multispectral optical and synthetic aperture radar (SAR) imagery (Shendryk, 2022; Sialelli et al., 2025; Nascetti et al., 2023; Pascarella et al., 2023). The emergence of remote sensing foundation models, which are large scale pretrained models that learn rich representations from satellite imagery (Feng et al., 2025; Zhu et al., 2024; Huang et al., 2025), provides an opportunity to improve feature quality for mapping tasks. These approaches typically use ensemble methods such as Random Forest (Breiman, 2001) or XGBoost (Chen and Guestrin, 2016) to predict biomass from satellite derived features at arbitrary locations.
@@
However, biomass point estimates alone are insufficient for decision making because uncertainty quantification is essential (Araza et al., 2022). Conservation planners need to know where estimates are reliable to prioritize field validation efforts and assess risk (Song et al., 2023). Carbon accounting requires quantifying uncertainty to meet reporting standards. Despite this need, most biomass mapping studies provide only limited uncertainty representations, typically reporting cross-validation RMSE or ensemble variance as proxies for predictive uncertainty (Li et al., 2020; Xu et al., 2024; Zurqani, 2025; Turton et al., 2022). To our knowledge, no prior work has systematically evaluated whether reported uncertainties for GEDI-based machine learning biomass mapping are actually calibrated, or proposed methods specifically designed to produce reliable prediction intervals. An important but often overlooked distinction exists between prediction intervals, which quantify uncertainty about individual observations, and ensemble variance, which measures disagreement between models fit on different data samples. When biomass mapping studies report uncertainty, do these uncertainties actually reflect the data?
@@
This gap is noteworthy for GEDI-based biomass mapping due to the mission’s irregular sampling pattern, and more broadly for ecological modeling where sparse but clustered sampling is the norm. Recent work has validated prediction interval coverage for GLM-based biomass estimation using airborne LiDAR (Mukhopadhyay et al., 2024), but equivalent scrutiny has not been applied to the ML ensemble methods standard in GEDI wall-to-wall mapping. Standard machine learning methods learn global relationships between satellite features and biomass, treating all predictions equivalently regardless of their spatial relationship to training data. Intuitively, however, uncertainty should depend on context. Predictions far from GEDI footprints should be more uncertain than those nearby, and predictions in heterogeneous landscapes should be more uncertain than in homogeneous areas. Ensemble based methods do not naturally encode these spatial relationships. Recent work has applied probabilistic deep learning to GEDI data using deep ensembles to estimate raw waveforms (Lang et al., 2022). However, their approach focuses on the 1D signal processing task (deriving height from a waveform at a single location). Our work addresses the 2D spatial interpolation task (predicting biomass between tracks), where uncertainty is driven by spatial distance and landscape heterogeneity.
In this work, we demonstrate that standard machine learning methods for biomass mapping face fundamental limitations in uncertainty quantification, and that these limitations persist across diverse forest biomes. The core issue is a conflation of two distinct statistical quantities. Ensemble variance, what Random Forest typically report as uncertainty, measures model disagreement across bootstrap samples or random initializations. Prediction intervals measure something different, namely the range within which future observations are likely to fall. These answer different questions, and the former systematically underestimates the latter when models are well-tuned for accuracy (Section 4.1). Compounding this, standard methods treat predictions as independent given input features, with no mechanism to adapt uncertainty to local spatial context. A prediction at a forest-field boundary and a prediction in contiguous forest receive identical uncertainty estimates if their spectral signatures match, despite vastly different local variability (Section 4.2).
@@
We introduce an alternative approach based on Neural Processes (NPs) (Garnelo et al., 2018; Kim et al., 2019), a class of models that explicitly condition predictions on observed data points and naturally produce calibrated uncertainty estimates. Geostatistics via Gaussian Processes (Rasmussen and Williams, 2005) provides the theoretically pure framework for this problem, but it scales cubically with the number of data points, making it operationally difficult for GEDI-scale datasets. Neural Processes offer a way to capture the spirit of GP-style spatial conditioning in a scalable, amortized fashion. We combine NPs with embeddings from a remote sensing foundation model trained on Sentinel-1 and Sentinel-2 data (Feng et al., 2025), demonstrating that this architecture can effectively map GEDI L4A biomass estimates to wall-to-wall predictions with well-calibrated uncertainty. Neural Processes treat any set of context observations $(x_{c},y_{c})$ as conditioning information, learn a distribution over functions given those contexts, and make predictions at arbitrary target locations $x_{t}$, which is directly aligned with the GEDI mapping problem.
@@
We provide a systematic evaluation of uncertainty calibration for GEDI-based biomass mapping, revealing that standard ensemble methods (Random Forest, XGBoost) exhibit poor calibration despite being widely used in the literature. Then, we show that Attentive Neural Processes achieve parity or better accuracy while maintaining near-ideal calibration. We validate these findings across five diverse regions spanning temperate (Maine, US), alpine (South Tyrol, Italy), boreal (Hokkaido, Japan), tropical lowlands (Guaviare, Colombia), and tropical montane (Tolima, Colombia) forests, demonstrating consistent performance across biomes. ANPs exhibit more graceful degradation under spatial extrapolation when trained on one region and tested on ecologically distinct areas. The model appropriately increases uncertainty, which is a useful property for operational deployment. Finally, we demonstrate that ANPs enable few-shot adaptation to new biomes, recovering functional accuracy with minimal local data, a capability that tree-based ensembles lack due to their non-differentiable structure.
@@
Our findings have implications for operational biomass mapping. The poor calibration of standard methods suggests that reported uncertainties may be unreliable, potentially leading to misguided conservation decisions or inaccurate carbon accounting. Neural Processes provide a principled alternative that produces trustworthy uncertainty estimates.
@@ 2.1.1 GEDI LIDAR Data
We obtain GEDI Level 4A (L4A) aboveground biomass density (AGBD) estimates using the gediDB library (Besnard et al., 2025), which provides efficient spatial queries. To facilitate alignment with satellite imagery and avoid spatial autocorrelation in validation, we organize L4A shots into a tile structure with 0.1 $\times$ 0.1 geographic tiles.
We apply quality filtering based on standard flags to exclude low confidence estimates (see appendix). Following standard practice for biomass modeling (Chave et al., 2014), we apply log transformation to stabilize variance and address the log-normal distribution of biomass data. Additionally, we filter out AGBD values over 500 Mg/ha following previous literature (Sialelli et al., 2025; Carreiras et al., 2017). Values exceeding this threshold (representing <1% of observations) are rare and likely reflect instrumental artifacts rather than true biomass, which are occasionally present in the GEDI L4A product even after quality filtering (e.g. physically unrealistic values exceeding 3000 Mg/ha).
We note that GEDI L4A AGBD values are themselves estimates with associated uncertainty (Dubayah et al., 2022). In this work, we treat GEDI estimates as the target variable. Consequently, the aleatoric uncertainty predicted by the ANP captures the conditional distribution of GEDI estimates given satellite features, incorporating both the variability of the landscape and the measurement uncertainty inherent in the GEDI product.
@@ 2.1.2 Remote Sensing Foundation Model Embeddings
Rather than using hand crafted spectral indices or raw satellite imagery, we use embeddings from Tessera (Feng et al., 2025), a transformer-based remote sensing foundation model pretrained on Sentinel-1 (C-band SAR) and Sentinel-2 (multispectral optical) imagery at global scale. We retrieve these embeddings via the geotessera Python package and use them as frozen features; Tessera itself is not fine-tuned. Tessera produces 128-dimensional embeddings at 10m spatial resolution that encode rich spectral-temporal-structural information about land surface characteristics.
@@ 2.2 Neural Process Architecture
We implement an Attentive Neural Process (ANPs) (Kim et al., 2019), which extend the Conditional Neural Process model (Garnelo et al., 2018) with attention mechanisms for improved context aggregation. Neural Processes are meta-learning models that learn distributions over functions by conditioning on a set of observed context points. Unlike standard supervised learning methods that learn a single function mapping inputs to outputs, NPs learn to produce a different predictive distribution for each set of context observations.
@@ 2.2.1 Architecture Components
Stochastic latent path: Mean pooled context representations parameterize a latent Gaussian distribution $\mathcal{N}(\mu_{z},\sigma_{z})$, which captures global uncertainty about the underlying function. Samples from this distribution (via the reparameterization trick (Kingma and Welling, 2014)) are combined with the deterministic path to form the final representation.
@@ 2.2.2 Training Objective
which encourages accurate predictions with appropriately scaled uncertainties. The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and prior, which regularizes the latent space. We use $\beta$-VAE warmup (Higgins et al., 2017) to prevent posterior collapse, linearly increasing $\beta$ over the first 10 epochs. This allows the model to first learn useful representations before regularization.
@@ 2.5 Baseline Models
Random Forest (RF): We use scikit-learn’s Random Forest implementation with ensemble sizes ranging from 50-1000 trees and maximum depths from 1-20. Uncertainty estimates are derived from the variance of predictions across trees as a simple baseline method. While computationally efficient and widely used, RF uncertainty estimates reflect ensemble variance (disagreement between trees) rather than true predictive uncertainty (Mentch and Hooker, 2016).
XGBoost: We use gradient boosting with quantile regression loss to estimate the 16th and 84th percentiles of the conditional distribution, approximating $\pm 1\sigma$ intervals assuming normality, alongside the mean estimate, as a more principled baseline of uncertainty quantification compared to ensemble variance. Hyperparameters range from tree depths 1-20 and boosting from 50-1000. Quantile regression provides explicit uncertainty estimates but fits global residual distributions.
@@ 2.7.1 Primary Study Region
Our primary experiments are conducted in the Guaviare department of Colombia (1 $\times$ 1 region from -73W to -72W, centered on 2.5N), which covers tropical forests experiencing varying degrees of disturbance, deforestation, and regeneration. This region provides diverse biomass conditions ranging from cleared land (0 Mg/ha) to intact forest (>200 Mg/ha), which allows model evaluation across the full range of interest for conservation and carbon monitoring applications.
@@ 3.4.2 Few-Shot Fine Tuning
Neural Processes are trained episodically. Each iteration samples a spatial tile, splits GEDI observations into context and target sets within that tile, and learns to predict targets conditioned on context. This episodic structure naturally enables few-shot learning, where the unit of adaptation is the tile, not individual GEDI footprints. We fine-tune source-region models on 10 tiles randomly sampled from the target region’s training split across 5 seeds, representing approximately 10-15% of the full training set. Models are trained for 5 epochs using the same optimizer settings and episodic procedure as initial training, continuing from the same 5 trained checkpoints for each region from Section 3.3.
@@
We do not present few-shot results for Random Forest or XGBoost because tree ensembles cannot be fine tuned in a similar way. Neural networks learn continuous, differentiable functions $f(x;\theta)$ where parameters $\theta$ can be adjusted via gradient descent. Tree ensembles consist of discrete decision rules (split thresholds, leaf values) that are fixed once determined and cannot be updated via gradients. The only options for ensemble methods given new data are to train from scratch on target-region data alone, incremental training by adding trees but new trees don’t adapt existing decisions, or combine source and target data and retrain fully (not few-shot learning, just more data with full retraining cost). This architectural limitation of discrete decision boundaries that cannot be smoothly adjusted prevents ensemble methods from leveraging pretraining for adaptation.

### v3

# Interpolation of GEDI Biomass Estimates with Calibrated Uncertainty Quantification

@@ Abstract
Reliable wall-to-wall biomass density estimation from NASA’s GEDI mission requires interpolating sparse LIDAR observations across heterogeneous landscapes. While machine learning approaches like Random Forest and XGBoost are widely used, they treat spatial predictions of GEDI observations from multispectral or SAR remote sensing data as independent without adapting to the varying difficulty of heterogeneous landscapes. We demonstrate these approaches generally fail to produce calibrated prediction intervals. We show that this stems from conflating ensemble variance with aleatoric uncertainty and ignoring local spatial context.
To resolve this, we introduce Attentive Neural Processes (ANPs), a probabilistic meta-learning architecture that explicitly conditions predictions on local observation sets and exploits geospatial foundation model embeddings. Unlike static ensembles, ANPs learn a flexible spatial covariance function, allowing estimates to be more uncertain in complex landscapes and less in homogeneous areas. We validate this approach across five distinct biomes ranging from tropical Amazonian forests to boreal, temperate, and alpine ecosystems, demonstrating that ANPs achieve competitive accuracy while maintaining near-ideal uncertainty calibration. We demonstrate the operational utility of the method through few-shot adaptation, where the model recovers most of the performance gap in cross-region transfer using minimal local data. This work provides a scalable, theoretically rigorous alternative to ensemble variance for continental scale earth observation.
@@ 1 Introduction
Forest aboveground biomass density (AGBD) is important for carbon accounting, climate change mitigation, and conservation planning (Dubayah et al., 2022; Duncanson et al., 2022). Accurate and high resolution biomass maps enable monitoring of deforestation, forest degradation, and carbon sequestration which are essential for implementing conservation programs and climate targets. NASA’s Global Ecosystem Dynamics Investigation (GEDI) mission provides high quality LIDAR measurements of forest structure, including calibrated estimates of AGBD at $\sim$25m footprint resolution (Dubayah et al., 2022). However, GEDI samples sparsely, with footprints separated by approximately 600m along-track in criss-crossing orbital patterns. To support operational forest monitoring, these sparse observations must be interpolated to create wall-to-wall maps.
@@
The emergence of remote sensing foundation models, which are large scale pretrained models that learn rich representations from satellite imagery (Feng et al., 2025; Zhu et al., 2024; Huang et al., 2025), provides an opportunity to improve feature quality for mapping tasks. Existing approaches to wall-to-wall biomass mapping from GEDI data typically aggregate shots to coarse resolution ($\sim$1km) to increase sampling density (Duncanson et al., 2022; Dubayah et al., 2022). Recently, machine learning and deep learning methods have enabled finer-resolution mapping (10-30m) by combining GEDI footprints with multispectral optical and synthetic aperture radar (SAR) imagery (Shendryk, 2022; Sialelli et al., 2025; Nascetti et al., 2023; Pascarella et al., 2023). These approaches typically use ensemble methods such as Random Forest (Breiman, 2001) or XGBoost (Chen and Guestrin, 2016) to predict biomass from satellite derived features at arbitrary locations.
@@
However, biomass point estimates alone are insufficient for decision making because uncertainty quantification is essential (Araza et al., 2022). Conservation planners need to know where estimates are reliable, to prioritize field validation efforts and assess risk (Song et al., 2023). Carbon accounting also requires quantifying uncertainty to meet reporting standards (International Organization for Standardization, 2018). Despite this need, most biomass mapping studies provide only limited uncertainty representations, typically reporting cross-validation RMSE or ensemble variance as proxies for predictive uncertainty (Li et al., 2020; Xu et al., 2024; Zurqani, 2025; Turton et al., 2022). To our knowledge, no prior work has systematically evaluated whether reported uncertainties for GEDI-based machine learning biomass mapping are actually calibrated (i.e. whether a 95% prediction interval actually contains the true value 95% of the time), or proposed methods specifically designed to produce reliable prediction intervals. An important but often overlooked distinction exists between prediction intervals, which quantify uncertainty about individual observations, and ensemble variance, which measures disagreement between models fit on different data samples. When biomass mapping studies report uncertainty, do these uncertainties actually reflect the data?
@@
This gap is noteworthy for GEDI-based biomass mapping due to the mission’s irregular sampling pattern, and more broadly for ecological modeling where sparse but clustered sampling is the norm. Recent work has validated prediction interval coverage for GLM-based biomass estimation using airborne LIDAR (Mukhopadhyay et al., 2024), but equivalent scrutiny has not been applied to the ML ensemble methods standard in GEDI wall-to-wall mapping. Standard machine learning methods learn global relationships between satellite features and biomass, treating all predictions equivalently regardless of their spatial relationship to training data. Intuitively, however, uncertainty should depend on context. Predictions for locations far from GEDI footprints should be more uncertain than those nearby, and predictions in heterogeneous landscapes should be more uncertain than in homogeneous areas. Ensemble based methods do not naturally encode these spatial relationships. Recent work has applied probabilistic deep learning to GEDI data using deep ensembles to estimate raw waveforms (Lang et al., 2022). However, their approach focuses on the 1D signal processing task (deriving height from a waveform at a single location). Our work addresses the 2D spatial interpolation task (predicting biomass between tracks), where uncertainty is driven by spatial distance and landscape heterogeneity.
In this work, we demonstrate that standard machine learning methods for biomass mapping face fundamental limitations in uncertainty quantification, and that these limitations persist across diverse forest biomes. We identify two core issues, namely the conflation of ensemble variance with prediction intervals, and inability to adapt uncertainty to local spatial context (Sections 4.14.2).
@@
We introduce an alternative approach based on Neural Processes (NPs) (Garnelo et al., 2018; Kim et al., 2019), a class of models that explicitly condition predictions on observed data points and naturally produce calibrated uncertainty estimates. Geostatistics via Gaussian Processes (Rasmussen and Williams, 2005) provides a theoretically sound framework for this problem, but it scales cubically with the number of data points, making it operationally difficult for GEDI-scale datasets. Neural Processes offer a way to capture the spirit of GP-style spatial conditioning in a scalable, amortized fashion. Neural Processes treat any set of context observations $(x_{c},y_{c})$ as conditioning information, learn a distribution over functions given those contexts, and make predictions at arbitrary target locations $x_{t}$, which is directly aligned with the GEDI mapping problem. We combine NPs with embeddings from a remote sensing foundation model trained on Sentinel-1 and Sentinel-2 data (Feng et al., 2025), demonstrating that this architecture can effectively map GEDI L4A biomass estimates to wall-to-wall predictions with well-calibrated uncertainty.
@@
We provide a systematic evaluation of uncertainty calibration for GEDI-based biomass mapping, revealing that standard ensemble methods (Random Forest, XGBoost) exhibit poor calibration. Then, we show that Attentive Neural Processes achieve parity or better accuracy while maintaining near-ideal calibration. We validate these findings across five diverse regions spanning temperate (Maine, US), alpine (South Tyrol, Italy), boreal (Hokkaido, Japan), tropical lowlands (Guaviare, Colombia), and tropical montane (Tolima, Colombia) forests, demonstrating consistent performance across biomes. ANPs exhibit more graceful degradation under spatial extrapolation when trained on one region and tested on ecologically distinct areas. The model increases uncertainty when necessary, a useful property for operational deployment. Finally, we demonstrate that ANPs enable few-shot adaptation to new biomes, recovering functional accuracy with minimal local data, a capability that decision tree-based ensembles lack.
@@
Our findings have implications for operational biomass mapping. The poor calibration of standard methods suggests that reported uncertainties may be unreliable, potentially leading to misguided conservation decisions or inaccurate carbon accounting. We provide a principled alternative that produces trustworthy uncertainty estimates.
@@ 2.1.1 GEDI LIDAR Data
We obtain GEDI Level 4A (L4A) aboveground biomass density (AGBD) estimates using the gediDB library (Besnard et al., 2025), which provides efficient spatial queries. To facilitate alignment with satellite imagery and avoid spatial autocorrelation during validation, we organize L4A shots into a tile structure with 0.1 $\times$ 0.1 geographic tiles.
We apply quality filtering based on standard flags to exclude low confidence estimates (see B). Following standard practice for biomass modeling (Chave et al., 2014), we apply log transformation to stabilize variance and address the log-normal distribution of biomass data. Additionally, we filter out AGBD values over 500 Mg/ha following previous literature (Sialelli et al., 2025; Carreiras et al., 2017). Values exceeding this threshold (representing <1% of observations) are rare and likely reflect instrumental artifacts rather than true biomass, which are occasionally present in the GEDI L4A product even after quality filtering. We note that GEDI L4A AGBD values are themselves estimates with associated uncertainty (Dubayah et al., 2022). In this work, we treat GEDI estimates as the target variable.
@@ 2.1.2 Remote Sensing Foundation Model Embeddings
Rather than using hand crafted spectral indices or raw satellite imagery, we use embeddings from Tessera (Feng et al., 2025), a transformer-based remote sensing foundation model pretrained on Sentinel-1 (C-band SAR) and Sentinel-2 (multispectral optical) imagery at global scale. We retrieve these embeddings via the geotessera Python package (Feng et al., 2025) and use them as frozen features; Tessera itself is not fine-tuned. Tessera produces 128-dimensional embeddings at 10m spatial resolution that encode rich spectral-temporal-structural information about land surface characteristics.
@@ 2.2 Neural Process Architecture
We implement an Attentive Neural Process (ANPs) (Kim et al., 2019), which extend the Conditional Neural Process model (Garnelo et al., 2018) with attention mechanisms for improved context aggregation. Neural Processes are meta-learning models that learn distributions over functions by conditioning on a set of observed context points. Unlike standard supervised learning methods that learn a single function mapping inputs to outputs, NPs learn to produce a different predictive distribution for each set of context observations. We now discuss ANPs in further detail, as these powerful tools are currently rare in the remote sensing literature.
@@ 2.2.1 Architecture Components
Stochastic latent path: Mean pooled context representations parameterize a latent distribution $\mathcal{N}(\mu_{z},\sigma_{z})$, which captures global uncertainty about the underlying function. Samples from this distribution (via the reparameterization trick (Kingma and Welling, 2014)) are combined with the deterministic path to form the final representation.
@@ 2.2.2 Training Objective
which encourages accurate predictions with appropriately scaled uncertainties. The second term is the Kullback-Leibler (KL) divergence between the approximate posterior and the learned prior, which regularizes the latent space. We use $\beta$-VAE warmup (Higgins et al., 2017) to prevent posterior collapse, linearly increasing $\beta$ over the first 10 epochs. This allows the model to first learn useful representations before regularization.
@@ 2.5 Baseline Models
Random Forest (RF): We use scikit-learn’s Random Forest implementation with ensemble sizes ranging from 50-1000 trees and maximum depths from 1-20. Uncertainty estimates are derived from the variance of predictions across trees as a simple baseline method. While computationally efficient and widely used, RF uncertainty estimates only reflect ensemble variance (disagreement between trees) rather than true predictive uncertainty (Mentch and Hooker, 2016).
XGBoost: We use gradient boosting with quantile regression loss to estimate the 16th and 84th percentiles of the conditional distribution, approximating $\pm 1\sigma$ intervals assuming normality, alongside the mean estimate, as a more principled baseline of uncertainty quantification compared to ensemble variance (Koenker and Bassett, 1978). Hyperparameters range from tree depths 1-20 and boosting from 50-1000. Quantile regression provides explicit uncertainty estimates but fits global residual distributions.
@@ 2.7.1 Primary Study Region
Our primary experiments are conducted in the Guaviare department of Colombia (1 $\times$ 1 region from -73W to -72W, centered on 2.5N), which covers tropical forests experiencing varying degrees of disturbance, deforestation, and regeneration. This region provides diverse biomass conditions ranging from cleared land to intact forest, which allows model evaluation across the full range of interest for conservation and carbon monitoring applications.
@@ 3.4.2 Few-Shot Fine Tuning
Recall that Neural Processes are trained episodically. Each iteration samples a spatial tile, splits GEDI observations into context and target sets within that tile, and learns to predict targets conditioned on context. This episodic structure naturally enables few-shot learning, where the unit of adaptation is the tile, not individual GEDI footprints. We fine-tune source-region models on 10 tiles randomly sampled from the target region’s training split across 5 seeds, representing approximately 10-15% of the full training set. Models are trained for 5 epochs using the same optimizer settings and episodic procedure as initial training, continuing from the same 5 trained checkpoints for each region from Section 3.3.
@@
We do not present few-shot results for Random Forest or XGBoost because tree ensembles cannot be fine tuned in this way. Neural networks learn continuous, differentiable functions $f(x;\theta)$ where parameters $\theta$ can be adjusted via gradient descent. Tree ensembles consist of discrete decision rules (split thresholds, leaf values) that are fixed once determined and cannot be updated via gradients. The only options for ensemble methods given new data are to train from scratch on target-region data alone, incremental training by adding trees–but new trees don’t adapt existing decisions–or combine source and target data and retrain fully (which is not few-shot learning, just more data with full retraining cost). This architectural limitation of discrete decision boundaries that cannot be smoothly adjusted prevents ensemble methods from leveraging pretraining for adaptation.